RFBG Commissioning with Adaptive Secondary - 01 April 2010 UT
Observer: JHill (LBTO)
Telescope Operator: WWack (until 03:00 UT) then JHill
Instrument Support:: JMorris
Telescope Support: JLittle
Adaptive Support: RBriguglio (LBTO)
Day crew fixed the leak in DX mirror ventilation.
JMH updated the DXPMIRTCFCollimation.dat file so that the M1 collimation uses the measured focus (Z) curve derived from the data collected on 20100329.
There were problems with the X-Y portion of the collimation data, so those are still using the shape derived from LBC-Red - as for the last 12 nights.
Andrew proposes that we should apply a "-15 arcsec Tilt" correction to M2 with the PSF arbitrator, and then take another set of field aberration data. Later he reconsidered and
would rather understand some of the other mysteries of DX collimation before doing that.
Plan for tonight (or whenever it clears): First, remap the residual Z6 on M2 as a function of elevation to recheck the lookup table compensation. It did not seem to be working well
at the end of last night. In good seeing, we should collect more transformation data. In moderate seeing, we should repeat the collection of collimation data (anticipating the M1 limits at
The second half of April 1, 2 UT are scheduled for W-unit testing.
Sky is mostly clear with scattered cirrus at sunset, but we are closed for high winds at the beginning of the night. The forecast is calling for rain, and then light snow later in the night.
Meanwhile, Runa has been able to set the shell at all elevations between 90 and 20 degrees.
We spent some of this closed-dome time making tests of shell performance as a function of telescope elevation.
By midnight winds are gusting above 30 m/sec.
Later Marco and Alfio were testing W-unit software and its coordination with telescope motion.
There was a power outage just after I went to bed at 10:30 UT.
- 31 Mar 2010